OpenStudy (cookiemonster18):

The table tell whether y varies directly with X if it does write an equation for the direct variation.

OpenStudy (cookiemonster18):

x-3,7,12 y-5.4,12.6,21.6 A.Not a direct variation B.direct variation,y=1.8x C.direct variation,y=4x D.direct variation,y=5/9x

OpenStudy (lochana):

answer is B

OpenStudy (lochana):

first try to plug each x and see if y values come up correctly. so x= 3 y = 1.8*3 = 5.4 others go same

OpenStudy (cookiemonster18):

can you help with 1 more @lochana

OpenStudy (lochana):

sure

OpenStudy (cookiemonster18):

The table tell whether y varies directly with X if it does write an equation for the direct variation. x=-2,3,8 y=1,6,11 A.Not a direct variation B.direct variation,y=x C.direct variation,y=2x D.direct variation,y=-2x

OpenStudy (anonymous):

it would be a